OpenStudy (anonymous):

If segment AB = 12" and sement BC = 8", then what is the value of 2(AB − BC) + 3(AB + 2BC)? 4(12) + 5(28) 2(4) + 3(28) 5(4) + 4(18) 5(2) − 4(18)

OpenStudy (anonymous):

sorry it is 2(12-8) + 3(12+2(8))

OpenStudy (anonymous):

i stil, couldnt fing the answer..

OpenStudy (anonymous):

12-8 = ? 12 + 2*8 = ?

OpenStudy (anonymous):

4, 28

OpenStudy (anonymous):

so 2(12-8)+3(12+2(8))=?

OpenStudy (anonymous):

84..

OpenStudy (anonymous):

no i meant 2(12-8)+3(12+2(8))=2(4)+3(28) = option 2

OpenStudy (anonymous):

2(4) + 3(28) ??

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you, your the best (:

OpenStudy (anonymous):

lol thank you. and you're welcome